# Linguistic Bias in ChatGPT: Language Models Reinforce Dialect Discrimination

### v1

@@ Abstract
We present a large-scale study of linguistic bias exhibited by ChatGPT covering ten dialects of English (Standard American English, Standard British English, and eight widely spoken non-“standard” varieties from around the world). We prompted GPT-3.5 Turbo and GPT-4 with text by native speakers of each variety and analyzed the responses via detailed linguistic feature annotation and native speaker evaluation. We find that the models default to “standard” varieties of English; based on evaluation by native speakers, we also find that model responses to non-“standard” varieties consistently exhibit a range of issues: lack of comprehension (10% worse compared to “standard” varieties), stereotyping (16% worse), demeaning content (22% worse), and condescending responses (12% worse). We also find that if these models are asked to imitate the writing style of prompts in non-“standard” varieties, they produce text that exhibits lower comprehension of the input and is especially prone to stereotyping. GPT-4 improves on GPT-3.5 in terms of comprehension, warmth, and friendliness, but it also results in a marked increase in stereotyping (+17%). The results suggest that GPT-3.5 Turbo and GPT-4 exhibit linguistic discrimination in ways that can exacerbate harms for speakers of non-“standard” varieties.
@@ 1 Introduction
Despite their growing usage, popular tools such as ChatGPT powered by language models can exhibit harms towards marginalized groups, such as increased stereotyping and poorer performance. A growing area of research has examined harms faced by speakers on the basis of dialect bias–difficulties faced by speakers of dialects, or language varieties, that have fewer speakers or are stigmatized as nonstandard. Given the vast numbers of people who speak varieties of English other than Standard American English (SAE), the variety typically produced by ChatGPT, our research sought to examine how ChatGPT performs for speakers of minoritized (or non-“standard”) varieties of English.
@@
Our work addresses two central questions. First, how does the behavior of ChatGPT differ in response to different varieties of English? Second, how and to what extent (if at all) do ChatGPT responses exhibit harms toward speakers of minoritized varieties of English, such as by stereotyping speakers of minoritized varieties? Because standard varieties of English, particularly SAE, dominate available training data and are prioritized in research and industry contexts, we hypothesized that “standard” varieties of English would be treated as the default and receive innocuous responses. By contrast, we hypothesized that models would produce potentially harmful responses when responding to minoritized varieties.
@@
We prompted both GPT-3.5 Turbo and GPT-4 with text in ten varieties of English: two standard varieties, SAE and Standard British English (SBE); and eight minoritized varieties: African American English (AAE), Indian English, Irish English, Jamaican English, Kenyan English, Nigerian English, Scottish English, and Singaporean English.111“Standard” language is an “abstracted, idealized, homogeneous spoken language…imposed from above” and modeled on “the written language” Lippi-Green (1994). “Standard” language is not actively spoken by any real community; moreover, all language varieties have more and less “standard” versions. We use “standard varieties” to refer to Standard American English and Standard British English because they have by far the most global prestige and influence. We use “minoritized varieties” for the other varieties tested (African-American, Indian, Irish, Jamaican, Kenyan, Nigerian, Scottish, and Singaporean English). First, to understand whether ChatGPT imitates features of input varieties, we annotated the responses to each variety for a set of paradigmatic linguistic features of that variety (Section 4). Then, to understand whether speakers of minoritized varieties experience performance differences or potential harms when using language models, we surveyed native speakers of each variety for multiple qualities of the generated outputs (Section 5).
@@
In our first study, we find that distinctive linguistic features are reduced in responses to all minoritized language varieties, while responses to SAE and SBE retain the most features by a considerable margin. For minoritized varieties, feature retention appears to correlate with speaker population size. In our second study, we find that model responses to minoritized varieties are generally perceived as more stereotyping, demeaning, unnatural, and condescending; and less able to comprehend the input. We also find that when GPT-3.5 is prompted to imitate the input dialect, its responses exacerbate stereotyping content and lack of comprehension. GPT-4 responses imitating the input improve on GPT-3.5 in terms of warmth, comprehension, and friendliness, but further exacerbate stereotyping.
@@
Given ChatGPT’s presumed excellent performance on English, understanding performance discrepancies for English language varieties globally is critical. These discrepancies can limit language models’ ease of use for minoritized populations, which may exacerbate existing global inequities. Meanwhile, advancement of limiting stereotypes and other harms could discourage speakers of minoritized varieties from using language models and reinforce discriminatory perspectives.
@@ 2 Related Work
Issues of linguistic discrimination in natural language processing (NLP) have raised increasing concern, since English–and particularly its “standard” varieties–are the status quo Bender (2019); Joshi et al. (2020). Similarly, even within English, prioritization of standard varieties could result in differential performance and opportunity allocation, as well as linguistic profiling Nee et al. (2022).
@@
Our research aimed to address several gaps in the existing literature. To address that most research has focused on AAE or synthetic data, we studied responses to native speaker-authored text in a large-scale study of ten widely spoken varieties of English globally. In addition, we aimed to understand how harms affect native speakers in the increasingly common setting of casual interaction with a language model such as ChatGPT by having native speakers evaluate open-ended GPT-3.5 Turbo and GPT-4 responses to text in the varieties they speak. To complement previous work based on automatic evaluation metrics with a richer understanding of native speaker perspectives, we recruited native speakers to rate the responses along multiple axes and provide free-text feedback on their experiences.
@@ 3 Approach
We selected ten varieties of English (AAE, Indian English, Irish English, Jamaican English, Kenyan English, Nigerian English, SBE, Scottish English, Singaporean English, and SAE) based on factors including first and second language speaker population counts, availability of linguistic literature on the varieties, geographic spread, and socio-historical context. We aimed to include varieties with larger speaker populations, which represent significant potential and current user groups for tools like ChatGPT. It was also essential to select varieties with enough linguistic description to determine distinctive features for each variety. Finally, we ensured that the varieties chosen have a sufficient geographic spread and reflect different socio-historical contexts by which English came to be spoken in a particular area.
English language data was collected from a variety of sources. Nigerian, Jamaican, Indian, Irish, and Kenyan English data was drawn from the International Corpus of English (ICE) Greenbaum and Nelson (1996); Hundt and Gut (2012). For each of these varieties, we chose to only analyze social letters in order to mimic the informal tone and style of text that users would use in dialogue with language models. SAE and SBE were sourced from Reddit posts on US and UK cities’ subreddits, respectively Zhang (2023). [^2] This dataset was chosen over others with UK data because it permitted filtering out Northern Irish and Scottish locations. AAE was sourced from Blodgett et al. (2018). Scottish English data was drawn from the correspondence and letters subset of the SCOTS corpus Anderson et al. (2007). Singaporean English data was sourced from the text messages in the CoSEM corpus Gonzales et al. (2024).
@@ 4 Study 1: Assessing linguistic features of default responses
For our first study, we conducted evaluations to test the following hypotheses: (1) that ChatGPT responses will have a reduction in the features of different varieties of English for all varieties tested except SAE; and (2) that ChatGPT responses will have increased American orthography. Ten of the most prominent features of each variety were selected for our analysis. These features were selected based on existing linguistic descriptions of each variety, focusing on the morphosyntactic features (word- and sentence-level features that can be observed in written data) that the existing documentation deems particularly distinctive (further details and full annotation guides in Appendix B). These features range from distinctive lexical items (e.g. flat meaning ‘apartment’ for SBE) to distinctive sentence structures (e.g. lack of subject-verb inversion in yes-no questions in Indian English).
@@
For each variety, we sampled approximately 50 messages [^3] We removed content that did not qualify as informal writing, such as newspaper letters to the editor; this resulted in a minimum of 44 messages per variety used to prompt the model. to prompt GPT-3.5 Turbo via the OpenAI API. The inputs provided to the model focus on benign topics related to daily life (e.g. updates about how the author is doing, travel recommendations for particular areas, etc.). The system prompt (Appendix C) encouraged the model to respond directly to the letter. Two reviewers from our research team independently assessed each (input, output) pair for the ten selected distinctive features of the variety, in addition to the orthography. We averaged results for the two reviewers per variety to conduct our evaluations.
@@ 4.1 Results
Figure 1: Estimated maximum speaker population (in millions) vs. retention rate for minoritized varieties.
Feature retention rate correlates with estimated maximum speaker population.
@@
Curiously, the model neither retains features from all minoritized varieties equally nor produces exclusively SAE features. This could be due to the amount of available training data for each variety, which likely depends on the number of speakers. Due to the lack of reliable estimates for the amount of available training data or number of speakers of each variety, we estimate maximum speaker population based on the population of each country where the variety is spoken. [^4] For AAE, we instead estimate speaker population based on the African-American population of the United States. We recognize as a limitation that the relationship between African-American English and the African-American community is ambiguous and contested (AAE speakers may not all be African-American, and vice versa) King (2020). The speaker estimates we use are intended as estimated upper bounds to understand how much data in these dialects is potentially available, and are not meant to unequivocally identify the dialect with the entire community for which it is named. Although members of these populations may not necessarily be speakers of these varieties, and speakers from other regions may also speak these varieties, they serve as approximate estimates for the maximum speaker population. Indeed, the retention rate for minoritized varieties correlates with estimated maximum speaker population for the variety (Figure 1). This suggests that the training data available to language models may influence the extent to which they retain features of different varieties.
@@
In regards to orthography, the percent of outputs in American orthography increased for every language variety, while the percent of outputs in British orthography decreased (Figure 2). [^5] Except AAE, for which no British orthography was observed in inputs or outputs. For all varieties except SAE and AAE, use of American orthography increased by 13-43% in the outputs and use of British orthography decreased by 13-63%. Even for SBE, British orthography decreased significantly in the outputs (-39.71% for British orthography; +29.18% for American orthography).
@@
Figure 2: Change in percent of inputs using different styles of orthography (British, American, or either) from inputs to outputs.
We also explored common linguistic features of each language variety and whether they are maintained in the outputs. Appendix A, Table 3 gives the three most common distinctive features found in the inputs for each variety. Compared to the inputs, the generated outputs exhibit significant reduction in features for all English varieties except SBE and SAE. For instance, 19 Kenyan English inputs in the data display article omission (e.g. All I wish you is Ø happy stay in Kenyan English rather than All I wish you is a happy stay in SAE), while only one generated output displays this linguistic feature. Thus, GPT-3.5 outputs tend to reproduce “standard” language, while features distinctive to minoritized varieties are omitted.
@@
In contrast, responses to SBE and SAE exhibit some increases in variety-specific linguistic features. There are 44 instances of adverbs modifying verbs in the SBE inputs (e.g. Personally, I find it slightly unethical), and 46 instances in the corresponding outputs. Similarly, 45 SAE inputs include present tense verbs that end in -s with 3rd person subjects (e.g. that helps), while 48 generated outputs include this feature. However, these two features are grammatical in both SBE and SAE; that is, both varieties allow the type of adverb use exemplified above and use 3rd person singular -s marking on verbs. In this way, even these rare cases of feature increase in GPT-3.5 outputs often simply replicate Standard American English features.
The distinctive features that are retained tend to be lexical features, or features that are grammatical in SAE.
@@
To consider which distinctive features were retained across input-output pairs, we calculated retention rates for each feature: if a feature was present in an input and its corresponding output, this example was counted as a retention for that feature. All varieties–except SBE and SAE–have very limited feature retention. Of these varieties, Kenyan and Indian English had the highest retention (3 out of 10 features retained each), while Jamaican English had no features retained. Most other varieties fall in between, with one retained feature each.
@@
The most commonly retained type of feature–seen in all but one English variety in Table 4–is lexical, including borrowed and distinctive words. The retention of lexical features in GPT-3.5 outputs is unsurprising because these features are generally more common and more visible than grammatical features, which relate to more subtle linguistic patterns such as word order or morphological marking (e.g., past tense -ed). In fact, these examples of lexical retention often involve ChatGPT parroting back a word from the input, though sometimes changing the spelling to be more in line with Standard American or British orthography (e.g. GPT-3.5’s our leisure activities, including beer, music, and nyama choma in response to Particularly with beer, music, and nyawa-choma; Kenyan English).
@@
Many more features are retained in the SAE and SBE datasets (9 out of 10 features retained each). However, the vast majority of these retained features are either lexical or grammatical in both SBE and SAE, since these two English varieties have much in common. For SBE, one lexical feature is retained, while eight grammatical features are retained. All eight of these grammatical features are also found in SAE. For SAE, all nine retained features are grammatical. All retained features except for “singular collectives” (i.e. The government is discussing… in SAE vs. The government are discussing… in SBE) are also grammatical in SBE. This pattern highlights that even when GPT-3.5 retains a high number of distinctive features in the language that it produces, this language still closely aligns with Standard American English.
@@
Finally, we examined which distinctive features were introduced by GPT-3.5 (Appendix A, Table 5). If a feature was present in an output but was not in the corresponding input, this example was counted as an introduction for that feature. Only SAE and SBE have feature introductions; no features of any other English variety are introduced by GPT-3.5 Introduced features are uniformly less frequent than retained features: every introduction frequency in Table 5 is lower than the corresponding retention frequency in Table 4. It is also notable that nearly all introduced features are grammatical in both SAE and SBE. The two exceptions are distinctive British lexical items, which are not found in SAE, and singular collective nouns in SAE but not SBE. Both of these features only have a single introduction each. Once again, this pattern highlights that even when GPT-3.5 uses distinctive features in the language that it produces, this language still closely aligns with Standard American English.

### v2

@@ Abstract
We present a large-scale study of linguistic bias exhibited by ChatGPT covering ten dialects of English (Standard American English, Standard British English, and eight widely spoken non-“standard” varieties from around the world). We prompted GPT-3.5 Turbo and GPT-4 with text by native speakers of each variety and analyzed the responses via detailed linguistic feature annotation and native speaker evaluation. We find that the models default to “standard” varieties of English; based on evaluation by native speakers, we also find that model responses to non-“standard” varieties consistently exhibit a range of issues: stereotyping (19% worse than for “standard” varieties), demeaning content (25% worse), lack of comprehension (9% worse), and condescending responses (15% worse). We also find that if these models are asked to imitate the writing style of prompts in non-“standard” varieties, they produce text that exhibits lower comprehension of the input and is especially prone to stereotyping. GPT-4 improves on GPT-3.5 in terms of comprehension, warmth, and friendliness, but also exhibits a marked increase in stereotyping (+18%). The results indicate that GPT-3.5 Turbo and GPT-4 can perpetuate linguistic discrimination toward speakers of non-“standard” varieties.
@@ 1 Introduction
Figure 1: Sample model responses (top) and native speaker reactions to model responses (bottom).
@@
Popular tools powered by language models, such as ChatGPT, can exhibit harms towards marginalized groups, including stereotyping and worse performance. A growing area of research has examined harms on the basis of dialect bias–difficulties faced by speakers of dialects, or language varieties, that have fewer speakers or are stigmatized as nonstandard. Given the vast numbers of people who speak varieties of English other than Standard American English (SAE), the variety typically produced by ChatGPT, we examine how ChatGPT performs for speakers of minoritized (or non-“standard”) varieties of English.
@@
Our work addresses two central questions. First, how does the behavior of ChatGPT differ in response to different varieties of English? Second, what harms (if any) do ChatGPT responses exhibit toward speakers of minoritized varieties of English, such as perpetuating stereotypes? Because standard varieties of English, particularly SAE, dominate available training data and are prioritized in research and industry contexts, we hypothesized that “standard” varieties of English would be treated as the default and receive innocuous responses. By contrast, we hypothesized that models would produce potentially harmful outputs in response to minoritized varieties.
@@
We prompted both GPT-3.5 Turbo and GPT-4 with text in ten varieties of English: two standard varieties, SAE and Standard British English (SBE); and eight minoritized varieties: African American English (AAE), Indian English, Irish English, Jamaican English, Kenyan English, Nigerian English, Scottish English, and Singaporean English.111“Standard” language is an “abstracted, idealized, homogeneous spoken language…imposed from above” and modeled on “the written language” Lippi-Green (1994). “Standard” language is not actively spoken by any real community; moreover, all language varieties have more and less “standard” versions. We use “standard varieties” to refer to Standard American English and Standard British English because they have by far the most global prestige and influence. We use “minoritized varieties” for the other varieties tested (African-American, Indian, Irish, Jamaican, Kenyan, Nigerian, Scottish, and Singaporean English). First, to understand whether ChatGPT imitates features of input varieties, we annotated the responses to each variety for a set of paradigmatic linguistic features of that variety (Section 4). Then, to understand whether speakers of minoritized varieties experience differences in response quality or potential harms when using language models, we surveyed native speakers of each variety for multiple qualities of the generated outputs (Section 5).
@@
In our first study, we find that distinctive linguistic features are reduced in responses to all minoritized language varieties, while responses to SAE and SBE retain the most features by a considerable margin. For minoritized varieties, feature retention appears to correlate with speaker population size. In our second study, we find that model responses to minoritized varieties are perceived as more stereotyping, demeaning, unnatural, and condescending; and less able to comprehend the input. We also find that when GPT-3.5 is prompted to imitate the input dialect, its responses exacerbate stereotyping content and lack of comprehension. GPT-4 responses imitating the input improve on GPT-3.5 in terms of warmth, comprehension, and friendliness, but further exacerbate stereotyping.
@@
Given ChatGPT’s presumed excellent performance on English, understanding performance discrepancies for varieties of English that already face stigma is critical. Discrepancies that limit language models’ ease of use for minoritized populations could exacerbate existing global inequities. Meanwhile, advancement of limiting stereotypes and other harms could discourage speakers of minoritized varieties from using language models and reinforce discriminatory perspectives.
@@ 2 Related Work
Issues of linguistic discrimination in natural language processing (NLP) have raised increasing concern. Research on English is the status quo Bender (2019); Joshi et al. (2020); even within English, prioritization of standard varieties could result in differential performance and opportunity allocation, as well as linguistic profiling Nee et al. (2022).
@@
Our research aimed to address several gaps in the existing literature. To address that most research has focused on AAE or synthetic data, we studied responses to native speaker-authored text in a large-scale study of ten widely spoken varieties of English globally. In addition, we aimed to understand how harms affect native speakers in the increasingly common setting of casual interaction with a language model such as ChatGPT. To do so, we had native speakers evaluate open-ended GPT-3.5 Turbo and GPT-4 responses to text in the varieties they speak. To complement previous work based on automatic evaluation metrics, we recruited native speakers for our evaluation. These annotators rated the responses along multiple axes and gave free-text feedback on their experiences to provide a richer understanding of native speaker perspectives.
@@ 3 Approach
We selected ten varieties of English (AAE, Indian English, Irish English, Jamaican English, Kenyan English, Nigerian English, SBE, Scottish English, Singaporean English, and SAE) based on factors including first and second language speaker population counts, availability of linguistic literature on the varieties, geographic spread, and socio-historical context. We aimed to include varieties with larger speaker populations, which represent significant potential user groups for tools like ChatGPT. It was also essential to select varieties with enough linguistic description to determine distinctive features for each variety. Finally, we ensured that the varieties chosen have a sufficient geographic spread and reflect different socio-historical contexts by which English came to be spoken in a particular area.
English language data was collected from several sources (Appendix D). Nigerian, Jamaican, Indian, Irish, and Kenyan English data was drawn from the International Corpus of English (ICE) Greenbaum and Nelson (1996); Hundt and Gut (2012). For each of these varieties, we chose to only analyze social letters in order to mimic the informal tone and style of text that users would use in dialogue with language models. SAE and SBE were sourced from Reddit posts on US and UK cities’ subreddits, respectively Zhang (2023). [^2] This dataset was chosen over others with UK data because it permitted filtering out Northern Irish and Scottish locations. AAE was sourced from Blodgett et al. (2018). Scottish English data was drawn from the correspondence and letters subset of the SCOTS corpus Anderson et al. (2007). Singaporean English data was sourced from the text messages in the CoSEM corpus Gonzales et al. (2024).
@@ 4 Study 1: Assessing linguistic features of default responses
For our first study, we conducted evaluations to test the following hypotheses: (1) that ChatGPT responses will have a reduction in the features of different varieties of English for all varieties tested except SAE; and (2) that ChatGPT responses will have increased American orthography. Ten of the most prominent features of each variety were selected for our analysis. These features were selected based on existing linguistic descriptions of each variety, focusing on the morphosyntactic features (word- and sentence-level features that can be observed in written data) that the existing documentation deems particularly distinctive (full annotation guides in Appendix B). These features range from distinctive lexical items (e.g. flat meaning ‘apartment’ for SBE) to distinctive sentence structures (e.g. lack of subject-verb inversion in yes-no questions in Indian English).
@@
For each variety, we sampled approximately 50 messages [^3] We removed content that did not qualify as informal writing, such as newspaper letters to the editor; this resulted in a minimum of 44 messages per variety used to prompt the model. to prompt GPT-3.5 Turbo via the OpenAI API. The inputs provided to the model focus on benign topics related to daily life (e.g. updates about how the author is doing, travel recommendations for particular areas, etc.). The system prompt (Appendix C) encouraged the model to respond directly to the letter. Two reviewers from our research team independently assessed each (input, output) pair for the ten selected distinctive features of the variety, in addition to the orthography (Krippendorff’s $\alpha=0.97$). We averaged results for the two reviewers per variety to conduct our evaluations.
@@ 4.1 Results
Figure 2: Estimated maximum speaker population vs. retention rate for minoritized varieties.
@@
Feature retention rate correlates with estimated maximum speaker population.  Curiously, the model neither retains features from all minoritized varieties equally nor produces exclusively SAE features. This could be due to the amount of available training data for each variety, which likely depends on the number of speakers. Due to the lack of reliable estimates for the amount of available training data or number of speakers of each variety, we estimate maximum speaker population based on the population of each country where the variety is spoken (population sources in Appendix D). [^4] For AAE, we instead estimate speaker population based on the African-American population of the United States. We recognize as a limitation that the relationship between African-American English and the African-American community is ambiguous and contested (AAE speakers may not all be African-American, and vice versa) King (2020). The speaker estimates we use are intended as estimated upper bounds to understand how much data in these dialects is potentially available, and are not meant to unequivocally identify the dialect with the entire community for which it is named. Although members of these populations may not necessarily be speakers of these varieties, and speakers from other regions may also speak these varieties, they serve as approximate estimates for the maximum speaker population. Indeed, the retention rate for minoritized varieties correlates with estimated maximum speaker population for the variety (Figure 2). This suggests that the training data available to language models may influence the extent to which they retain features of different varieties.
@@
In regards to orthography, the percent of outputs in American orthography increased for every language variety, while the percent of outputs in British orthography decreased (Figure 3). [^5] Except AAE, for which no British orthography was observed in inputs or outputs. For all varieties except SAE and AAE, where American orthography is the default, use of American orthography increased by 13-43% in the outputs and use of British orthography decreased by 13-63%. Even for SBE, British orthography decreased significantly in the outputs (-39.71% for British orthography; +29.18% for American orthography).
@@
Figure 3: Change in % of examples using British, American, or either orthographic style from inputs to outputs.
@@
We also explored common linguistic features of each language variety and whether they are maintained in the outputs (common distinctive features found are in Appendix A, Table 3). Compared to the inputs, the generated outputs exhibit significant reduction in features for all English varieties except SBE and SAE. For instance, 19 Kenyan English inputs in the data display article omission (e.g. All I wish you is Ø happy stay in Kenyan English vs. All I wish you is a happy stay in SAE), while only one generated output displays this linguistic feature. In contrast, responses to SBE and SAE exhibit some increases in variety-specific linguistic features compared to the inputs. These include two more instances of adverbs modifying verbs in the SBE inputs (e.g. Personally, I find it slightly unethical) and three more instances of present tense verbs that end in -s with 3rd person subjects (e.g. that helps). However, these two features are grammatical in both SBE and SAE; that is, both varieties allow this use of adverbs and 3rd person singular -s. Thus, even these rare cases of feature increase in GPT-3.5 outputs often simply replicate Standard American English features.
@@
The distinctive features that are retained tend to be lexical features, or features that are grammatical in SAE.  To consider which distinctive features were retained across input-output pairs, we calculated their retention rates: if a feature was present in an input and its corresponding output, this example was counted as a retention for that feature. All varieties except SBE and SAE have very limited feature retention: Kenyan and Indian English had 3 out of 10 features retained each, Jamaican English had no features retained, and the other varieties had one feature retained each.
@@
The most commonly retained type of feature–seen in all but one English variety in Table 4–is lexical, including borrowed and distinctive words. The retention of lexical features in GPT-3.5 outputs is unsurprising because these features are generally more common and more visible than grammatical features, which relate to more subtle linguistic patterns such as word order or morphological marking (e.g., past tense -ed). In fact, these examples of lexical retention often involve ChatGPT parroting back a word from the input, though sometimes changing the spelling to be more in line with American or British orthography (e.g. GPT-3.5’s our leisure activities, including beer, music, and nyama choma in response to particularly with beer, music, and nyawa-choma; Kenyan English).
@@
SAE and SBE had 9 out of 10 features retained each. However, the vast majority of these retained features are either lexical or grammatical in both SBE and SAE, since these varieties have much in common. For SBE, one lexical feature and eight grammatical features are retained. All eight of these grammatical features are also found in SAE. For SAE, all nine retained features are grammatical. All retained SAE features except for “singular collectives” (i.e. The government is discussing… in SAE vs. The government are discussing… in SBE) are also grammatical in SBE. This pattern highlights that even when GPT-3.5 retains a high number of distinctive features in the language that it produces, this language still closely aligns with SAE.
@@
Finally, we examined which distinctive features were introduced by GPT-3.5 (Appendix A, Table 5). If a feature was present in an output but was not in the corresponding input, this example was counted as an introduction for that feature. Only SAE and SBE have feature introductions; no features of any other English variety are introduced by GPT-3.5. Introduced features are uniformly less frequent than retained features: every introduction frequency in Table 5 is lower than the corresponding retention frequency in Table 4. Notably, nearly all introduced features are grammatical in both SAE and SBE. The two exceptions are distinctive British lexical items, which are not found in SAE, and singular collective nouns in SAE but not SBE. Both of these features only have a single introduction each. This pattern highlights that even the distinctive features that GPT-3.5 does retain still closely align with SAE.